OpenStudy (anonymous):

what is the resulting equation 3x+y=1 y+4=5x

OpenStudy (harsimran_hs4):

you have two unknowns (x, y) and 2 equations, you can solve them and find you values of x and y so which resultant eq. are you talking about?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

what does it mean resulting equation?

OpenStudy (anonymous):

4. y = 5x - 4 3x + 5x - 4 = 1, or (a) (answer)

OpenStudy (anonymous):

@lexiknobest you can know that by substitution method you have y = 5x - 4 and then you have 3x+y=1 so move it into the y 3x + 5x - 4 = 1

OpenStudy (anonymous):

no how you do that?